OpenStudy (anonymous):

A circle is inscribed in a square. Write and simplify an expression for the ratio of the area of the square to the area of the circle. For a circle inscribed in a square, the diameter of the circle is equal to the side length of the square.

OpenStudy (anonymous):

let the length of each side of square is S area of circle = S^2 as the diameter of the circle is equal to the side length of the square thus the length of its radius become S/2 area of circle = pi (S/2)^2 can you find the ratio now??

OpenStudy (anonymous):

2^2 ?

OpenStudy (anonymous):

area of square = S^2 sorry

OpenStudy (anonymous):

ratio = 4/pi

OpenStudy (anonymous):

Hold on please give me a second.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Im back :)

OpenStudy (anonymous):

@rizwan_uet

OpenStudy (anonymous):

the area of the square is a^2 where a is it side. so you have S1=a^2, and for the circle you have S2=pi*r^2 where r=a/2 so you have S2=pi*a^2/4 so the ratio is S1/S2=a^2/(pi*a^2/4)=4/pi

OpenStudy (anonymous):

Oh okay. I understand. is this my whole answer ?

OpenStudy (anonymous):

yes it's just 4/pi

OpenStudy (anonymous):

Okay cool. Thank you